OpenStudy (credmond):

Can somone help me with my lab? It is a M1V1=M2V2 equation. Just need help setting it up. We labeled six 50.00mL volumetric flasks 1-6. Then pipet 10.00 mL of .200 M Fe(NO3)3 solution into each volumetric flask. Then pipet 1.00, 2.00, 3.00, 4.00, 5.00 mL of 2.00 x 10^-3 M NaSCN solution into flasks 2 through 6 respectively. We then added sufficient 0.10 M nitric acid to each flask to make the total volume 50.00 mL. The equation we are to use is M1V2=M2V2 to find the initial [SCN-], M. If you can help me set it up for each flask number that would be great! :) Thanks!

OpenStudy (aaronq):

\( Fe^{3+}+SCN^- \rightleftharpoons [Fe(SCN)]^{2+}\) ; formation constant: \(K_f\)=? so find how many moles of SCN \((n_{SCN^−})\) are left after the complexation? HINT: write an equilibrium expression Your final volume is 50 mL so 0.05 L with unknown molarity "x M" \(xM=\dfrac{n_{SCN^−}}{L_{solution}}→xM=\dfrac{n_{SCN^−}}{0.05L}\)

OpenStudy (credmond):

Oh, okay gotcha. So for the first equation would it be (2x10^-3)(1mL)/.05L?

OpenStudy (aaronq):

not quite. you have to find how much SCN is left after the complexation of \([Fe(SCN)]^{2+}\) occurs.

OpenStudy (credmond):

How do I find that?

OpenStudy (aaronq):

write an equilibrium expression (make an ICE table) and find how much of it is left after everything has equilibrated.

OpenStudy (credmond):

Would you mind doing the first equation for me? I should be able to understand it after that.

OpenStudy (aaronq):

the equilibrium expression is: \(K_f=\dfrac{[Fe(SCN)^{2+}]}{[Fe^{3+}][SCN^-]}\) okay, at 25 degrees celsius \(K_f^{[Fe(SCN)]^{2+}}=8.9×10^2\) \(8.9×10^2=\dfrac{[Fe(SCN)^{2+}]}{[Fe^{3+}][SCN^-]}\) "Then pipet 10.00 mL of .200 M Fe(NO3)3 solution into each volumetric flask" \(n_{Fe^{3+}}=L_{Fe(NO_3)_3}*M_{Fe(NO_3)_3}=(0.001\;L)(0.200 M)=0.0002\;moles\) new molarity: \(M_{Fe^{3+}}=\dfrac{0.0002\;moles}{0.05L}=0.004 M\) flask 2: "Then pipet 1.00 mL o f 2.00 x 10^-3 M NaSCN" (through a similar calc. to the one above) \(M_{SCN^-}=0.00004\;M\) \(Fe^{3+} + SCN^- \rightleftharpoons [Fe(SCN)]^{2+}\) I 0.004 0.00004 0 C -x -x +x E 0.004-x 0.00004-x x \(8.9×10^2=\dfrac{x}{(0.004-x)\color{red}{(0.00004-x)}}\rightarrow x=0.0000311743708865\) \([SCN^-]=\color{red}{(0.00004-x)}=(0.00004-0.0000311743708865)\) \([SCN^-]=0.000008825629=8.8*10^{-6}\) for flask 2.

OpenStudy (aaronq):

by the way i didn't mention that i substituted the values from E (from the ICE table) into the initial equilibrium expression \(K_f=\dfrac{[Fe{SCN)]^{2+}}}{[Fe^{3+}][SCN^-]}\)